OpenStudy (anonymous):

Integral question, seemed easy for me, but I crushed into a big wall. Please help me. (I'll post the question)

OpenStudy (anonymous):

\[\Large \int\limits_{-1}^{1}~e^{u+1}~~du\]I am rewriting it below, to just integrate e^u.\[\Large \int\limits_{-1}^{1}~e(e)^{u}~~du\]The indefinite integral,\[\Large \frac{e(e)^{u+1}}{u+1}\] And I get an undefined value for the lower sum-:(

OpenStudy (anonymous):

The integral of \(e^u\) is just \(e^u\), not \(\dfrac{e^{u+1}}{u+1}\).

OpenStudy (anonymous):

The power rule only applies to constant exponents, not variable ones.

OpenStudy (anonymous):

ohh damn, I suck so much, tnx

OpenStudy (anonymous):

I myself adjusted it to e^u, and then say this bs. Sorry...

OpenStudy (anonymous):

I won't do these silly mistakes anymore. I'll pay more attention. Tnx again, S&G

OpenStudy (anonymous):

You're welcome!